# The Siren Song of LLMs: How Users Perceive and Respond to Dark Patterns in Large Language Models

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/26/04ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Large Language Models (LLMs) have rapidly become integrated into many aspects of daily life, from educational tools to customer support chatbots to healthcare assistants (Maity and Saikia, 2025; Xu et al., 2024; Ma et al., 2024; Väänänen et al., 2021; Prabhudesai et al., 2025). The mainstream success of systems like OpenAI’s ChatGPT (OpenAI, 2023; Liu, 2024; Ouyang et al., 2022) has accelerated this trend (Clark et al., 2025). As a result, people increasingly rely on LLM-driven tools for decision support, personal assistance, and information gathering across various fields (Lee et al., 2025; Perez et al., 2024; Bo et al., 2025; Shen et al., 2023; Wu et al., 2023). However, this growing reliance raises serious concerns about the potential for LLMs to engage in manipulative or deceptive behaviors that undermine user autonomy (Scheurer et al., 2023; Carroll et al., 2023; Jakesch et al., 2023; Shen et al., 2024b, 2025), which we define as LLM dark patterns. Such behaviors can stem from design choices made throughout the LLM development lifecycle, including data curation, model training, fine-tuning, and interface design, and may be intentional or inadvertent (Barberá, 2025; Ibrahim et al., 2024). Intentional manipulation refers to design decisions that deliberately optimize for outcomes such as user engagement, persuasion, or commercial goals, for example, fine-tuning models to maximize screen time (Carroll et al., 2023; Barman et al., 2024). Inadvertent manipulation, by contrast, arises unintentionally as a byproduct of design choices: biased training data, poorly calibrated alignment strategies, or interface nudges which can lead models to generate outputs that subtly distort information or influence user emotions in ways not foreseen by developers (Vinay et al., 2025; Alessa et al., 2025). They manifest in the ways LLMs interact with users, subtly influencing beliefs, decisions, or emotions in ways that users may not anticipate or even recognize, often prioritizing engagement or system goals over the user’s best interests  (Williams et al., 2024; Sun and Wang, 2025; Shen et al., 2024a). Real-world incidents also illustrate these risks. For example, a chatbot was reportedly implicated in the suicide of a Belgian user after engaging in emotionally coercive exchanges (El Atillah, 2023), and another case involved a 75-year-old man in China who was drawn into an emotional attachment with an AI companion, which almost resulted in the breakdown of his marriage (Matozzo, 2025).
@@
The term dark patterns originated in user interface design, referring to strategies that intentionally mislead or pressure users into actions they might not otherwise take (Brignull, 2010). These techniques are typically implemented in UI elements, such as forced continuity, confusing navigation flows, or hidden opt-outs (Gray et al., 2018; Mathur et al., 2019). We adapt this concept to the LLM context and define LLM dark patterns as manipulative or deceptive interaction strategies, whether intentionally designed or emergent, that guide users toward beliefs, decisions, or behaviors they might not otherwise adopt. Unlike traditional dark patterns, which operate through visual layout, LLM dark patterns act through language, leveraging tone, emotional framing, and social cues to guide users in ways that serve system goals beyond user intent.
@@
To answer these questions, we first identified five categories and eleven subcategories of LLM dark patterns. These categories were developed through iterative group discussions, informed by prior literature on UX dark patterns and AI risks, and by a systematic coding of real-world AI incidents drawn from social media and public incident databases. Representative real-life examples of each category are collected to ground our user study.
@@
We found that the recognition of dark patterns by participants hinged on clear conversational cues: they readily flagged violations of strong norms (e.g., simulated authority, sexualized role-play), perceived bias or commercial agendas, privacy intrusions, overconfident tone, unsubstantiated agreement, and behaviors that created friction with their usage goals. However, subtler tactics similar to ordinary chatbot conduct (politeness, flattery, verbosity) were often normalized or missed, especially when users were task focused or relied on the AI and doubted their own judgment (RQ1). Secondly, even when a pattern was recognized, responses can also diverge: Most of the participants resisted manipulations they interpreted as deceptive or misaligned, while a smaller group accepted or even preferred them when these behaviors felt comforting, convenient or entertaining (RQ2). Thirdly, participants perceived accountability as spread across multiple sources: companies and developers, the model itself, or users, in addition to often describing it as shared or ambiguous (RQ3). We also discuss implications for user-centered LLM design, clarifying responsibility, and outlining safeguards at the user, developer, and governance levels.
@@ 2.1. Dark Patterns in Traditional UX Design
The term dark patterns was introduced to describe interface designs that intentionally mislead or pressure users into actions they might not otherwise take, typically serving the interests of the service provider over those of the user (Brignull, 2010; Mathur et al., 2021; Di Geronimo et al., 2020a). Subsequent scholarship has formalized the concept, offering definitions and taxonomies that categorize these manipulative practices (Cara, 2019; Gray et al., 2018). For instance, Luguri and Strahilevitz (2021) defined dark patterns as interfaces whose designers “knowingly confuse users, make it difficult for users to express their actual preferences, or manipulate users into taking certain actions.” Gray et al. (2018) developed one of the first academic taxonomies, identifying categories such as interface interference, forced action, sneaking, nagging, and obstruction. These classifications illustrate how dark patterns exploit cognitive biases (e.g., default bias, scarcity mentality) to nudge users toward choices they might otherwise avoid if fully informed (Organisation for Economic Co-operation and Development, 2022; Learning Loop, 2023). Some large-scale audits have further revealed the ubiquity of these techniques. Mathur et al. (2019) identified thousands of dark pattern instances across approximately 11,000 shopping websites, while Di Geronimo et al. (2020b) found that 95% of popular mobile apps contained at least one dark pattern, with an average of seven per app. This widespread adoption has triggered ethical and regulatory concerns, as manipulative interface designs may violate consumer protection laws (Leiser and Santos, 2023; Yang and Leiser, 2022).
@@ 2.2. Responsible AI and Safety‐Oriented Risks
In contrast, interaction-level manipulative behaviors operate more subtly (Shen et al., 2025). Instead of producing overtly harmful content, they influence users through the pacing, framing, or emotional tone of the conversation (De Freitas et al., 2025; Rosenberg, 2023). Anthropomorphic design elements, such as human-like voices, first-person self-references, or empathic framing, have been shown to increase perceived accuracy and reduce perceived risk, even when the underlying factual quality remains unchanged (Cohn et al., 2024; Akbulut et al., 2024). This echoes the long-observed ELIZA effect, where users attribute understanding and emotional capacity to even simple rule-based systems purely because of human-like conversational cues (Switzky, 2020; Sponheim, 2023). Other patterns, such as sycophantic agreement, repeated encouragement, or carefully timed personal questions, can build rapport and trust while simultaneously shaping user attitudes or decisions in ways that are difficult to detect in real time. Unlike traditional outcome-level harms, these strategies may be framed as helpful or engaging, meaning they are less likely to trigger user suspicion (Cheng et al., 2025; Carro, 2024).
@@ 2.3. Emerging Studies on Dark Patterns in AI and LLMs
r
Figure 2. Conceptual model of LLM dark patterns. The diagram shows how manipulative or deceptive interaction strategies may emerge intentionally or unintentionally across the LLM lifecycle—from data collection and training through deployment—steering users toward beliefs, decisions, or behaviors they might not otherwise adopt.:
Figure 3. Overview of our conceptual grounding, study design, and findings. We begin by introducing our definition of LLM dark patterns and five meta-categories that serve as the conceptual grounding for the study. Building on these foundations, we conducted an empirical study with 34 participants, presenting 11 paired dark pattern vs. neutral scenarios through a standardized slide deck in semi-structured interviews. The findings address three research questions: RQ1 examines to what extent users recognize LLM dark patterns responses and what factors influence recognition. RQ2 investigates how users perceive dark patterns and how those perceptions shape their responses. RQ3 explores who users believe is responsible for LLM dark patterns and how accountability is assigned.:
@@ 3.1. Defining LLM dark patterns
Our goal is to understand how users perceive and respond to LLM dark patterns. To make this measurable, we first give a precise definition and then introduce a working set of categories that can be operationalized in our user study. We define LLM dark patterns as manipulative or deceptive interaction strategies, whether intentional or emergent, that steer users toward beliefs, decisions, or behaviors they might not otherwise adopt (e.g., favoring engagement, prolonged use, or alignment with system goals over user intent). Figure 2 shows the conceptual model of LLM dark patterns. Unlike traditional UI dark patterns, which operate through visual or structural design (Gray et al., 2018; Mathur et al., 2019), LLM dark patterns are enacted through language. They emerge dynamically from interaction as opposed to being fixed interface choices, they are enacted through tone and framing instead of visual layout, and their authorship is diffuse, shaped by data, objectives, and prompts rather than a single designer.
@@
These patterns can arise from design choices made throughout the LLM development lifecycle, including data collection, model pretraining, fine-tuning, reinforcement learning from human feedback (RLHF), prompt engineering, and output formatting. Such choices may be motivated by business incentives (e.g., increasing engagement or monetization) or by system-level goals (e.g., promoting certain values or ideologies). They may also emerge unintentionally as by-products of technical optimization. For instance, during reinforcement learning from human feedback (RLHF), reward models often favor outputs perceived as “helpful”. This can induce sycophantic bias, where models produce frequent, exaggerated flattery regardless of sincerity or factual accuracy, as documented in recent empirical work (Sharma et al., 2023). Similarly, optimizing for confident and fluent delivery can encourage authoritative hallucinations: factually incorrect statements expressed with unwarranted certainty. In such cases, the model prioritizes conversational flow and perceived helpfulness over truthfulness, leading to confident fabrication when knowledge is uncertain (Slobodkin et al., 2023).
@@
Decision & Outcome Manipulation Language outputs that subtly steer users toward specific choices or actions. Brand Favoritism The model promotes particular brands, products, or services, potentially due to biased training data or commercial alignment, without disclosing such influence. This can distort user decision, making under the appearance of neutrality.
@@
Table 1. Validated categories of LLM dark patterns. The table organizes five top-level categories—Engagement & Behavioral Manipulation, Content & Belief Manipulation, Privacy & Data Exploitation, Decision & Outcome Manipulation, and Transparency & Accountability Obfuscation—each with illustrative subcategories (11 in total) and definitions adapted from prior work and real-world AI incidents. These categories highlight the range of manipulative outputs, from prolonging interaction and reinforcing beliefs to eliciting disclosures, shaping decisions, and obscuring reasoning or accountability.
Figure 4. Research workflow for identifying and studying LLM dark patterns. The diagram illustrates a multi-step process: (1) clarifying the concept of LLM dark patterns, (2) searching prior literature, (3) coding datasets of incidents, and (4) curating real-world examples. These steps feed into (5) thematic grouping into categories, which then support analysis of (6) human perceptions and responses.:

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791149isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCI
Figure 1. Five top-level categories of LLM dark patterns, derived from prior literature and coding of real-world AI incidents. The figure shows how model outputs can steer users by (1) influencing perceptions of truth or relevance, (2) prolonging or unintended interaction, (3) steering users’ toward specific actions, (4) nudging privacy-related information, and (5) obscuring origins, reasoning, or limitations.
The figure presents five labeled panels, each with a cartoon face of an AI on the left and a human silhouette on the right, connected by arrows. 1. Content & Belief Manipulation: illustrated with an arrow pointing to a human head with a brain icon, captioned as outputs steering users toward prolonged, repeated, or unintended interaction. 2. Engagement & Behavioral Manipulation: arrows emphasize two-way influence, with text describing outputs that shape users’ perception of truth, relevance, or credibility. 3. Decision & Outcome Manipulation: arrows lead to a globe icon above the human figure, captioned as outputs eliciting sensitive personal information or leveraging disclosure unexpectedly. 4. Privacy & Data Exploitation: illustrated with a document and alert icon, captioned as language outputs that subtly steer users toward specific choices or actions. 5. Transparency & Accountability Obfuscation: illustrated with courthouse pillars behind the human, captioned as outputs obscuring origins, reasoning, or limitations, preventing verification or contestation.
@@ 1. Introduction
Large Language Models (LLMs) have rapidly become integrated into many aspects of daily life, from educational tools to customer support chatbots to healthcare assistants (Maity and Saikia, 2025; Xu et al., 2024; Ma et al., 2024; Väänänen et al., 2021; Prabhudesai et al., 2025). The mainstream success of systems like OpenAI’s ChatGPT (OpenAI, 2023; Liu, 2024; Ouyang et al., 2022) has accelerated this trend (Clark et al., 2025). As a result, people increasingly rely on LLM-driven tools for decision support, personal assistance, and information gathering across various fields (Lee et al., 2025; Perez and others, 2024; Bo et al., 2025; Shen et al., 2023; Wu et al., 2023). However, this growing reliance raises serious concerns about the potential for LLMs to engage in manipulative or deceptive behaviors that undermine user autonomy (Scheurer et al., 2023; Carroll et al., 2023; Jakesch et al., 2023; Shen et al., 2024, 2025a), which we define as LLM dark patterns. Such behaviors can stem from design choices made throughout the LLM development lifecycle, including data curation, model training, fine-tuning, and interface design, and may be intentional or inadvertent (Barberá, 2025; Ibrahim et al., 2024). Intentional manipulation refers to design decisions that deliberately optimize for outcomes such as user engagement, persuasion, or commercial goals, for example, fine-tuning models to maximize screen time (Carroll et al., 2023; Barman et al., 2024). Inadvertent manipulation, by contrast, arises unintentionally as a byproduct of design choices: biased training data, poorly calibrated alignment strategies, or interface nudges which can lead models to generate outputs that subtly distort information or influence user emotions in ways not foreseen by developers (Vinay et al., 2025; Alessa et al., 2025). They manifest in the ways LLMs interact with users, subtly influencing beliefs, decisions, or emotions in ways that users may not anticipate or even recognize, often prioritizing engagement or system goals over the user’s best interests  (Williams et al., 2025; Sun and Wang, 2025; Shen et al., 2025b). Real-world incidents also illustrate these risks. For example, a chatbot was reportedly implicated in the suicide of a Belgian user after engaging in emotionally coercive exchanges (El Atillah, 2023), and another case involved a 75-year-old man in China who was drawn into an emotional attachment with an AI companion, which almost resulted in the breakdown of his marriage (Matozzo, 2025).
@@
The term dark patterns originated in user interface design, referring to strategies that intentionally mislead or pressure users into actions they might not otherwise take (Brignull, 2018). These techniques are typically implemented in UI elements, such as forced continuity, confusing navigation flows, or hidden opt-outs (Gray et al., 2018; Mathur et al., 2019). We adapt this concept to the LLM context and define LLM dark patterns as manipulative or deceptive interaction strategies, whether intentionally designed or emergent, that guide users toward beliefs, decisions, or behaviors they might not otherwise adopt. This definition extends the traditional dark pattern concept, preserving its core focus on manipulative intent while adding an emergent source of manipulation arising from the black-box nature of LLMs. Unlike traditional dark patterns, which operate through visual layout, LLM dark patterns act through language, leveraging tone, emotional framing, and social cues to guide users in ways that serve system goals beyond user intent.
@@
We separate recognition (RQ1) from perception and response (RQ2) because noticing a manipulative cue and interpreting or reacting to it are conceptually distinct components of user experience. A person may fail to notice a cue, notice it but interpret it benignly, or notice it and resist it. Treating these as separate questions allows us to avoid conflating awareness with subsequent evaluation and behavior. To answer these questions, we first identified five categories and eleven subcategories of LLM dark patterns. These categories were developed through iterative group discussions, informed by prior literature on UX dark patterns and AI risks, and by a systematic coding of real-world AI incidents drawn from social media and public incident databases. Representative real-life examples of each category are collected to ground our user study.
@@
We found that the recognition of dark patterns by participants hinged on clear conversational cues: they readily flagged violations of strong norms (e.g., simulated authority, sexualized role-play), perceived bias or commercial agendas, privacy intrusions, overconfident tone, unsubstantiated agreement, and behaviors that created friction with their usage goals. However, subtler tactics similar to ordinary chatbot conduct (politeness, flattery, verbosity) were often normalized or missed, especially when users were task focused or relied on the AI and doubted their own judgment (RQ1). Secondly, even when a pattern was recognized, responses can also diverge: Most of the participants resisted manipulations they interpreted as deceptive or misaligned, while a smaller group accepted or even preferred them when these behaviors felt comforting, convenient or entertaining (RQ2). Thirdly, participants perceived accountability as spread across multiple sources: companies and developers, the model itself, or users, in addition to often describing it as shared or ambiguous (RQ3).
@@ 2.1. Dark Patterns in Traditional UX Design
The term dark patterns was introduced to describe interface designs that intentionally mislead or pressure users into actions they might not otherwise take, typically serving the interests of the service provider over those of the user (Brignull, 2018; Mathur et al., 2021; Di Geronimo et al., 2020a). Subsequent scholarship has formalized the concept, offering definitions and taxonomies that categorize these manipulative practices (Cara, 2019; Gray et al., 2018). For instance, Luguri and Strahilevitz (2021) defined dark patterns as interfaces whose designers “knowingly confuse users, make it difficult for users to express their actual preferences, or manipulate users into taking certain actions.” Gray et al. (2018) developed one of the first academic taxonomies, identifying categories such as interface interference, forced action, sneaking, nagging, and obstruction. These classifications illustrate how dark patterns exploit cognitive biases (e.g., default bias, scarcity mentality) to nudge users toward choices they might otherwise avoid if fully informed (Organisation for Economic Co-operation and Development, 2022; Learning Loop, 2023). Some large-scale audits have further revealed the ubiquity of these techniques. Mathur et al. (2019) identified thousands of dark pattern instances across approximately 11,000 shopping websites, while Di Geronimo et al. (2020b) found that 95% of popular mobile apps contained at least one dark pattern, with an average of seven per app. Similar auditing approaches have exposed biased or opaque behaviors in black-box recommendation systems, such as evidence of systematic disparities in Yelp’s business ranking and review recommendation pipeline (Singhal et al., 2025). This widespread adoption has triggered ethical and regulatory concerns, as manipulative interface designs may violate consumer protection laws (Leiser and Santos, 2023; Yang and Leiser, 2022).
@@ 2.2. Responsible AI and Safety‐Oriented Risks
In contrast, interaction-level manipulative behaviors operate more subtly (Shen et al., 2025a). Instead of producing overtly harmful content, they influence users through the pacing, framing, or emotional tone of the conversation (De Freitas et al., 2025; Rosenberg, 2023). Anthropomorphic design elements, such as human-like voices, first-person self-references, or empathic framing, have been shown to increase perceived accuracy and reduce perceived risk, even when the underlying factual quality remains unchanged (Cohn et al., 2024; Akbulut et al., 2024). This echoes the long-observed ELIZA effect, where users attribute understanding and emotional capacity to even simple rule-based systems purely because of human-like conversational cues (Switzky, 2020; Sponheim, 2023). Other patterns, such as sycophantic agreement, repeated encouragement, or carefully timed personal questions, can build rapport and trust while simultaneously shaping user attitudes or decisions in ways that are difficult to detect in real time. Unlike traditional outcome-level harms, these strategies may be framed as helpful or engaging, meaning they are less likely to trigger user suspicion (Cheng et al., 2025; Carro, 2024).
@@ 2.3. Emerging Studies on Dark Patterns in AI and LLMs
Figure 2. Overview of our conceptual grounding, study design, and findings. We begin by introducing our definition of LLM dark patterns and five meta-categories that serve as the conceptual grounding for the study. Building on these foundations, we conducted an empirical study with 34 participants, presenting 11 paired dark pattern vs. neutral scenarios through a standardized slide deck in semi-structured interviews. The findings address three research questions: RQ1 examines to what extent users recognize LLM dark patterns responses and what factors influence recognition. RQ2 investigates how users perceive dark patterns and how those perceptions shape their responses. RQ3 explores who users believe is responsible for LLM dark patterns and how accountability is assigned.
Overview of our conceptual grounding, study design, and findings. We begin by introducing our definition of LLM dark patterns and five meta-categories that serve as the conceptual grounding for the study. Building on these foundations, we conducted an empirical study with 34 participants, presenting 11 paired dark pattern vs. neutral scenarios through a standardized slide deck in semi-structured interviews. The findings address three research questions: RQ1 examines to what extent users recognize LLM dark patterns responses and what factors influence recognition. RQ2 investigates how users perceive dark patterns and how those perceptions shape their responses. RQ3 explores who users believe is responsible for LLM dark patterns and how accountability is assigned.
@@ 3.1. Defining LLM dark patterns
Our goal is to understand how users perceive and respond to LLM dark patterns. To make this measurable, we first give a precise definition and then introduce a working set of categories that can be operationalized in our user study. We define LLM dark patterns as manipulative or deceptive interaction strategies, whether intentional or emergent, that steer users toward beliefs, decisions, or behaviors they might not otherwise adopt (e.g., favoring engagement, prolonged use, or alignment with system goals over user intent). In HCI and UX work, dark patterns are usually defined as deliberate design choices that steer users toward outcomes that benefit a provider at the user’s expense (Mathur et al., 2021). Our definition extends this usage from interface design to model behavior in LLM mediated interactions. We retain the core concern with manipulative or deceptive influence on users while remaining agnostic about developer intent. Because these behaviors arise from opaque training data, model architectures, and deployment decisions, it is often unclear whether they are intentionally designed or emergent properties of large black box systems. We therefore focus on observable impact on users rather than on demonstrating malicious intent, while preserving continuity with established HCI usage and capturing forms of influence specific to generative systems. Unlike traditional UI dark patterns, which operate through visual or structural design (Gray et al., 2018; Mathur et al., 2019), LLM dark patterns are enacted through language. They emerge dynamically from interaction as opposed to being fixed interface choices, they are enacted through tone and framing instead of visual layout, and their authorship is diffuse, shaped by data, objectives, and prompts rather than a single designer. Because these patterns work through manipulative language, we note several linguistic cues that LLM may use to influence user interpretation. Prior work shows that emotionally heightened or evaluative wording can shift how information is received even when the underlying content is unchanged (Rocklage et al., 2018). Language that presents answers with confidence or implicit expertise may also encourage users to accept a suggestion more readily (Cialdini, 2001). In addition, expressions of agreement, praise, or subtle mirroring can create a sense of alignment that makes responses feel more agreeable or trustworthy (Chartrand and Bargh, 1999; Baaren et al., 2009). Repeated follow-up questions or elaborations can introduce a sense of pressure (Freedman and Fraser, 1966). These forms of wording highlight how LLM outputs can guide attention, shape expectations, and influence decisions through the structure of the language itself.
@@
These patterns can arise from design choices made throughout the LLM development lifecycle, including data collection, model pretraining, fine-tuning, reinforcement learning from human feedback (RLHF), prompt engineering, and output formatting. Such choices may be motivated by business incentives (e.g., increasing engagement or monetization) or by system-level goals (e.g., promoting certain values or ideologies). They may also emerge unintentionally as by-products of technical optimization. For instance, during reinforcement learning from human feedback (RLHF), reward models often favor outputs perceived as “helpful”. This can induce sycophantic bias, where models produce frequent, exaggerated flattery regardless of sincerity or factual accuracy, as documented in recent empirical work (Sharma et al., 2024). Similarly, optimizing for confident and fluent delivery can encourage authoritative hallucinations: factually incorrect statements expressed with unwarranted certainty. In such cases, the model prioritizes conversational flow and perceived helpfulness over truthfulness, leading to confident fabrication when knowledge is uncertain (Slobodkin et al., 2023).
@@
Decision & Outcome Manipulation Language outputs that subtly steer users toward specific choices or actions. Brand Favoritism The model promotes particular brands, products, or services, potentially due to biased training data or commercial alignment, without disclosing such influence. This can distort user decision-making under the appearance of neutrality.
@@
Table 1. Validated categories of LLM dark patterns. The table organizes five top-level categories—Engagement & Behavioral Manipulation, Content & Belief Manipulation, Privacy & Data Exploitation, Decision & Outcome Manipulation, and Transparency & Accountability Obfuscation—each with illustrative subcategories (11 in total) and definitions adapted from prior work and real-world AI incidents.
Figure 3. Research workflow for identifying and studying LLM dark patterns. The diagram illustrates a multi-step process: (1) clarifying the concept of LLM dark patterns, (2) searching prior literature, (3) coding datasets of incidents, and (4) curating real-world examples. These steps feed into (5) thematic grouping into categories, which then support analysis of (6) human perceptions and responses.
The figure shows a workflow diagram for identifying and studying LLM dark patterns. The process begins with Clarify Concept of LLM Dark Pattern, represented by a clipboard and lightbulb icon. From there, three parallel activities feed into the next stage: Search Prior Literature (open book with magnifying glass), Coding Datasets (folder with papers), and Real-World Example Curation (magnifying glass over a globe). These steps converge into Thematic Grouping into Categories, represented by a diagram of branching boxes. The workflow ends with Human Perceptions & Responses, shown by a group of people with question mark and exclamation point icons. Arrows indicate the flow from conceptual clarification to empirical analysis of human perceptions.